# Sample-efficient Unsupervised Policy Cloning from Ensemble Self-supervised Labeled Videos

### v1

@@ Abstract
Current advanced policy learning methodologies have demonstrated the ability to develop expert-level strategies when provided enough information. However, their requirements, including task-specific rewards, expert-labeled trajectories, and huge environmental interactions, can be expensive or even unavailable in many scenarios. In contrast, humans can efficiently acquire skills within a few trials and errors by imitating easily accessible internet video, in the absence of any other supervision. In this paper, we try to let machines replicate this efficient watching-and-learning process through Unsupervised Policy from Ensemble Self-supervised labeled Videos (UPESV), a novel framework to efficiently learn policies from videos without any other expert supervision. UPESV trains a video labeling model to infer the expert actions in expert videos, through several organically combined self-supervised tasks. Each task performs its own duties, and they together enable the model to make full use of both expert videos and reward-free interactions for advanced dynamics understanding and robust prediction. Simultaneously, UPESV clones a policy from the labeled expert videos, in turn collecting environmental interactions for self-supervised tasks. After a sample-efficient and unsupervised (i.e., reward-free) training process, an advanced video-imitated policy is obtained. Extensive experiments in sixteen challenging procedurally-generated environments demonstrate that the proposed UPESV achieves state-of-the-art few-shot policy learning (outperforming five current advanced baselines on 12/16 tasks) without exposure to any other supervision except videos. Detailed analysis is also provided, verifying the necessity of each self-supervised task employed in UPESV.
@@ I INTRODUCTION
With the advancement of reinforcement learning (RL), it’s possible to train policies that reach expert levels for challenging tasks defined in complex environments [1, 2, 3, 4]. However, the demanding training conditions and low sample efficiency significantly constrain the applicability of RL. To this end, many advanced RL approaches, such as model-based RL [5, 6, 7] and self-supervised RL [8, 3, 9], are proposed for higher sample efficiency. At the same time, some other researchers are no longer limited to online agent experience but try to seek help from offline supervisions. Given enough expert-labeled demonstrations, ideal policies can be obtained through offline-RL [10, 11] or imitation learning [12, 13], without the need of reward-specific environmental interactions. However, similar to expert rewards, these offline supervisions are also not widely available and always require time-consuming collection at a high cost. Different from them, the image-only videos are easily accessible currently thanks to the development of video websites and social media. However, Learning From Videos (LFV) poses a big challenge due to the lack of action labels. Although some attempts, such as video-based RL pre-training [14, 15] and video-based intrinsic rewards [16], have made great progress on RL data efficiency, they still require more or less additional supervisory information, e.g., reward-based finetuning and expert-annotated actions. Inverse RL methods [17, 18, 19] provide effective solutions for completely unsupervised policy learning from videos by reward prediction. While they often fail to well balance policy performance and sample efficiency, requiring more interactions than RL with expert rewards [20, 12].
@@
Figure 1: UPESV achieves sample-efficient policy learning with only action-free videos and reward-free interactions. This is achieved by three organically combined self-supervised tasks, where each performs its own duties. For example, on Fruitbot task, UPESV is the only effective few-shot policy learning method, where each self-supervised task is necessary.
@@
Can we derive a video-based policy with only a few interactions, without exposure to any other supervisions? Previous advanced methods [21, 22] have proven it possible and achieved considerable results in state-based control (with the state-based demonstrations) and style-unchanged easy video games. However, their performance is unsatisfactory when facing environments with complex visual input and dynamics, which may be attributed to the following two shortcomings: First, they don’t make full use of available information, i.e., both the expert videos and reward-free interactions. For example, BCO [21] tries to label expert videos only based on non-expert environmental interactions, which means the highly qualified dynamics information contained in expert videos and the distribution difference are ignored. In contrast, ILPO [22] pays much more attention to the above dynamics contained in videos, but their world model is trained without considering the potential assistance provided by ground-truth interactions. Second, they don’t provide their model with enough vision-specific signals. These signals are crucial for the models to understand high-dimensional inputs [23, 24], thus well extracting extensive information (e.g., environmental dynamics) from both datasets.
@@
In response to the above two issues, we propose a novel, sample-efficient, and unsupervised video-based policy learning framework, named Unsupervised Policy from Ensemble Self-supervised labeled Videos (UPESV), as shown in Fig. 1. It organically employs several self-supervised tasks to well utilize both expert videos and reward-free interactions, training a robust dynamics-aware video labeling model to infer the expert actions for efficient policy cloning. For expert videos, we employed a latent reconstruction task, where another world model is set to recover the next observation based on current observation and predicted actions, in the latent space. This task forces the labeling model to understand the changes between neighboring observations and extract the high-quality dynamics information contained in expert videos, which is highly related to expert action prediction. For reward-free interactions, we let the labeling model predict the non-expert ground-truth actions. This task not only aligns the model to the real action space, but also serves as a generalization task by changing both the data distribution and optimization target, thereby improving the model robustness. In practice, we observe a performance drop when naively gathering the above tasks into one model because they can’t learn a shared and qualified representation module by themselves. This motivates us to design a visual shift contrast task, where the model is required to associate a shifted image to its origin. This task let the model pay more attention to action-related visual changes, enhancing its representation ability in a target manner, which is proven necessary to well balance the above two tasks. The organically employed three self-supervised tasks together enable advanced labeling performance, where each performs its own duties. Along with the labeling model training, UPESV simultaneously clones a policy from the labeled expert video, and collects environmental interactions to enrich self-supervised task data in turn. After a sample-efficient, unsupervised, and iterative training process, UPESV obtains an advanced policy based on a robust video labeling model.
@@
We conduct extensive experiments in sixteen challenging procedurally-generated environments containing complex visual inputs. The results demonstrate that the proposed UPESV achieves considerable few-shot policy learning, outperforming current advanced baselines on 12/16 tasks without exposure to any other supervisions except videos. We also provide detailed ablations and analyses to show the necessity of all three self-supervised tasks employed in our framework. We summarize the contributions as follows:
@@
  • We propose a novel, sample-efficient, and unsupervised framework to derive policies from videos, named Unsupervised Policy from Ensemble Self-supervised labeled Videos (UPESV). Given a small amount of environmental interactions, UPESV can derive an effective video-based strategy, without need of any other supervision.
@@
  • UPESV trains a video labeling model along with imitating a policy, through three organically combined self-supervised tasks where each performs its own duties. These tasks together provide the models with advanced representation learning and comprehensive dynamics understanding.
@@
  • Extensive experiments in sixteen challenging procedurally-generated environments demonstrate that UPESV exhibits state-of-the-art performance in few-shot policy learning. It outperforms five current advanced baselines on 12/16 tasks. Detailed analyses are also provided to verify the effectiveness and necessity of each module in our framework design.
@@ II-A Sample-efficient Policy Learning
Interacting with the environment to collect data for policy learning is a time-consuming and expensive process. In some special scenarios, such as autonomous driving [25] and real robot learning [26], the interactions can even lead to danger. To this end, many advanced methods are proposed to improve the RL sample efficiency. Model-based RL methods [6, 7, 27, 28] train extra world models that augment the RL experience, thereby improving sampling efficiency[29]. Self-supervised RL, i.e., RL auxiliary tasks [30, 8, 31, 32] alleviate the lack of supervision signals for representation learning in DRL, accelerating policy learning by improving the upstream representation understanding capabilities. At the same time, many researchers try to completely avoid accessing environments, learning policies through imitation learning [12, 33] and offline RL [11, 10, 14]. However, the required expert supervision signals, including both expert reward and offline expert dataset, are also expensive and unavailable in many scenarios. Different from these expert supervisions, videos are easy to obtain from the internet currently. Our work aims to achieve completely unsupervised policy learning with the easily accessible videos, in the absence of any other expert supervision.
Figure 2: UPESV framework. UPESV learns a video labeling model and a policy network jointly, through three organically combined self-supervised tasks where each is necessary and performs its own duties. The motivation and details of these three tasks: visual shift contrast, latent future reconstruction, and ground-truth action prediction, are provided separately in Sections III.A, III.B, and III.C. We simultaneously train the policy $\pi(a|o)$ by behavior cloning labeled expert videos for self-supervised data enrichment, which we detail in Section III.D.
@@ II-B Learning Policies From Videos
Thanks to the improvement of the internet, it’s very easy to acquire massive and different kinds of videos currently. Many videos contain expert demonstrations in different fields, which provides extensive knowledge for both human beings and machines to utilize. However, due to the lack of action information, how to well utilize the videos poses a big challenge. One of the most direct ways is to improve existing strategy learning methods with the help of videos, such as video-based pre-training for RL [34, 35, 36, 37] and video-based intrinsic reward for RL [16]. They indeed improve the sample efficiency of policy learning under enough supervision but still require other expert supervision. Imitation Learning from only Observations (ILO) [17, 21, 22] is proposed to completely decouple expert supervisions from LFV. Most ILO methods try to extract expert rewards contained in the videos, employing RL to optimize the unsupervised reward expectation for policy imitation [17, 38, 39]. These inverse RL methods achieve considerable results on imitation performance but can’t take sample efficiency into account. At the same time, some researchers give up reward engineering for RL. By predicting expert actions from videos based on non-expert interactions [21, 40] or digging out expert dynamics hidden in the expert videos [22, 20], they achieve sample-efficiency unsupervised policy learning faced with simple inputs. Different from these methods, UPESV organically employs multiple different unsupervised tasks, benefiting from both non-expert interactions and expert videos with advanced representation learning.
@@ III Methodology
UPESV trains a policy network and a video labeling model jointly. The policy determines the action $a$ based on current observation $o$, defined as $\pi(a|o)$. The video labeling model predicts the real expert action between two neighboring expert observations in videos, formulated as a inverse dynamics model $V(a^{v}_{i}|o^{v}_{i},o^{v}_{i+1})$, where $a^{v}_{i}$ denotes the predicted action while $o^{v}_{i}$ and $o^{v}_{i+1}$ denote neighboring expert observations. $V$ consists of three networks: a feature encoder $f$, a latent predictor $g$, and an action projector $h$. $o^{v}_{i}$ and $o^{v}_{i+1}$ are separately encoded by $f$, concatenated, and processed by $g$ and $h$ to obtain the $a^{v}_{t}$, formulated as:
$a^{v}_{t}=h(g(f(o^{v}_{i}),f(o^{v}_{i+1}))).$ (1)
In practice, we also include a few available historical observations $o^{v}_{h}$ as additional input information. They are processed similar to $o^{v}_{i}$ and we omit them in the remaining part for clarity. UPESV learns $V$ through three different but related self-supervised tasks over both expert videos $B^{v}$ and environmental interactions $B^{e}$, which we describe in the first three subsections. During the training of $V$, we simultaneously train the policy $\pi(a|o)$ by behavior cloning $V$-labeled expert videos, which we describe in Section III.D. The policy $\pi(a|o)$ collects self-supervised data for $V$ while the improved $V$ further enhances the policy $\pi(a|o)$. After a sample-efficient and iterative training process, we obtain the well-trained video labeling model $V$ and policy $\pi$.
We provide the framework diagram, as shown in Fig. 2.
@@ III-A1 Motivation
To infer actions from videos, the video labeling model $V(a^{v}_{i}|o^{v}_{i},o^{v}_{i+1})$ should be able to capture the difference between neighboring images. In most visual control domains, the relative difference is more related to decision making than the absolute difference in visual inputs. Imagine two particles moving upward at different positions on the plane. They have different absolute positions, but the same relative position changes that corresponds to their same actions. To this end, we want the labeling model to focus on relative position difference, which is achieved by driving it to ignore the absolute position difference. Therefore, a shift matching task that forces the video labeling model to align a random shifted image to its origin in the semantic space, is employed. This task enhances the visual understanding of the $V$ in a target manner, which is necessary to understand complex dynamics through another two tasks (detailed in the next two subsections) simultaneously.
@@ III-A2 Implementation
The visual shift contrast task forces the video labeling model $V$ to align a randomly shifted image to its origin in the latent semantic space. Concretely, a batch of observations $\{o^{v}_{i}\}_{i=1}^{N}$ are sampled from expert video dataset $B^{v}$. Each observation is randomly shifted (up to $s$ units in any direction) twice to obtain two shifted images, $\hat{o}^{v}_{i}$ and $\check{o}^{v}_{i}$, where the two different hats upon ${o}^{v}_{i}$ denote two independent random shifting. They are further encoded by $f$ to obtain two latent features, $f(\hat{o}^{v}_{i})$ and $f(\check{o}^{v}_{i})$. Then, we define ($f(\hat{o}^{v}_{i})$,$f(\check{o}^{v}_{i})$) as positive pairs while ($f(\hat{o}^{v}_{i})$,$f(\hat{o}^{v}_{j})$) as negative pairs. Contrastive learning [41] is conducted on the whole batch to bring the positive pair closer while pulling the negative pair further away from each other. A contrastive model measuring distance is defined as follows:
$M_{c}(f(\hat{o}^{v}_{i}),f(\check{o}^{v}_{i}))=W\frac{u(f(\hat{o}^{v}_{i}))\cdot f(\check{o}^{v}_{i})}{||u(f(\hat{o}^{v}_{i}))||\cdot||f(\check{o}^{v}_{i})||},$ (2)
where $W$ is a trainable matrix. $u(\cdot)$ is a Multi-Layer Perceptron (MLP), which doesn’t change the dimension of the latent embedding. It is set to introduce asymmetry, thereby avoiding collapse to trivial solutions. Then, we define a InfoNCE loss [42] as the objective of visual shift contrast:
$\mathcal{L}_{VSC}=$ $\log M_{c}(f(\hat{o}^{v}_{i}),f(\check{o}^{v}_{i}))$ (3)
$-\log\frac{1}{N}\sum_{j=1}^{N}M_{c}(f(\hat{o}^{v}_{i}),f(\hat{o}^{v}_{j})).$

### v2

@@ Abstract
Current advanced policy learning methodologies have demonstrated the ability to develop expert-level strategies when provided enough information. However, their requirements, including task-specific rewards, action-labeled expert trajectories, and huge environmental interactions, can be expensive or even unavailable in many scenarios. In contrast, humans can efficiently acquire skills within a few trials and errors by imitating easily accessible internet videos, in the absence of any other supervision. In this paper, we try to let machines replicate this efficient watching-and-learning process through Unsupervised Policy from Ensemble Self-supervised labeled Videos (UPESV), a novel framework to efficiently learn policies from action-free videos without rewards and any other expert supervision. UPESV trains a video labeling model to infer the expert actions in expert videos through several organically combined self-supervised tasks. Each task performs its duties, and they together enable the model to make full use of both action-free videos and reward-free interactions for robust dynamics understanding and advanced action prediction. Simultaneously, UPESV clones a policy from the labeled expert videos, in turn collecting environmental interactions for self-supervised tasks. After a sample-efficient, unsupervised, and iterative training process, UPESV obtains an advanced policy based on a robust video labeling model. Extensive experiments in sixteen challenging procedurally generated environments demonstrate that the proposed UPESV achieves state-of-the-art interaction-limited policy learning performance (outperforming five current advanced baselines on 12/16 tasks) without exposure to any other supervision except for videos.
@@ I INTRODUCTION
With the advancement of reinforcement learning (RL) [1, 2, 3, 4], it’s possible to train policies that reach expert levels for challenging tasks defined in complex environments [5, 6, 7, 8]. However, the demanding training conditions and low sample efficiency significantly constrain the applicability of RL. To this end, many advanced RL approaches, such as model-based RL [9, 10, 11, 12] and self-supervised RL [13, 14, 15, 16], are proposed for higher sample efficiency. At the same time, some other researchers are no longer limited to online agent experience but try to seek help from offline supervisions. Given enough expert-labeled demonstrations, ideal policies can be obtained through offline RL [17, 18] or imitation learning [19, 20], without the need for reward-specific environmental interactions. However, similar to expert rewards, these offline supervisions are also not widely available and always require time-consuming collection at a high cost. Different from them, the image-only videos are easily accessible currently thanks to the development of video websites and social media. However, Learning From Videos (LFV) poses a big challenge due to the lack of action labels. Although some attempts, such as video-based RL pre-training [21, 22] and video-based intrinsic rewards [23], have made great progress on RL data efficiency, they still require more or less additional supervisory information, e.g., reward-based finetuning and expert-annotated actions. Inverse RL methods [24, 25, 26] provide effective solutions for completely unsupervised policy learning from videos by reward prediction. While they often fail to well balance policy performance and sample efficiency, requiring more interactions than RL with expert rewards [27, 19].
@@
Figure 1: UPESV achieves sample-efficient policy learning with only action-free videos and reward-free interactions. This is achieved by three organically combined self-supervised tasks, where each performs its duties. For example, on the interaction-limited Fruitbot task, UPESV obtains the only effective policy, where each self-supervised task is necessary.
@@
Can we derive a video-based policy with only a few interactions, without exposure to any other supervision? Previous advanced methods [28, 29] have proven it possible and achieved considerable results in state-based control (with the state-based demonstrations) and style-unchanged easy video games. However, their performance is unsatisfactory when facing environments with complex visual input and dynamics, which may be attributed to the following issues: They don’t make full use of available information, i.e., both the action-free expert videos and reward-free interactions. For example, BCO [28] tries to label expert videos only based on non-expert environmental interactions, which means the highly qualified dynamics information contained in expert videos and the distribution difference are ignored. In contrast, ILPO [29] pays much more attention to the above dynamics contained in videos, but their world model is trained without considering the potential assistance provided by ground-truth interactions. In addition, their models are not provided with enough representation-specific signals. These signals are crucial for the models to well understand complex high-dimensional inputs [30, 31, 32], thus extracting extensive action-related information (e.g., environmental dynamics) from both datasets.
@@
In response to the above issues, we propose a novel, sample-efficient, and unsupervised video-based policy learning framework, named Unsupervised Policy from Ensemble Self-supervised labeled Videos (UPESV), as shown in Fig. 1. It organically employs several self-supervised tasks to well utilize both action-free expert videos and reward-free interactions, training a robust and dynamics-aware video labeling model to infer the expert actions for efficient policy cloning. For action-free videos, we employed a latent reconstruction task, where another world model is set to recover the next observation based on the current observation and predicted actions, in the latent space. This task forces the labeling model to understand the changes between neighboring observations and extract the high-quality dynamics information contained in expert videos, which is highly related to expert action prediction. For reward-free interactions, we let the labeling model predict the non-expert ground-truth actions. This task not only aligns the model to the real action space but also serves as a generalization task by changing both the data distribution and optimization target, thereby improving the model’s robustness. In practice, we observe that naively gathering the above tasks into our framework can’t yield effective policies in some environments, which can be mainly attributed to the unqualified representation learning. This motivates us to design a visual shift contrast task, where the representation module is required to associate two shifted images with the same origin. This task let the video labeling model pay more attention to action-related visual changes, enhancing its visual understanding ability in a target manner to well support the above two tasks. The organically employed three self-supervised tasks together enable advanced labeling performance, where each performs its duties. Along with the labeling model training, UPESV simultaneously clones a policy from the labeled expert videos, and collects environmental interactions to enrich self-supervised task data in turn. After a sample-efficient, unsupervised, and iterative training process, UPESV obtains an advanced policy based on a robust video labeling model.
@@
We conduct extensive experiments in sixteen challenging procedurally generated environments with complex visual inputs. Results demonstrate that the proposed UPESV achieves state-of-the-art policy learning under limited interactions, outperforming current advanced baselines on 12/16 tasks without exposure to any other supervision except for videos.
We summarize the contributions as follows:
@@
  • We propose a novel, sample-efficient, and reward-free framework to learn from action-free videos, named Unsupervised Policy from Ensemble Self-supervised labeled Videos (UPESV). Given a few environmental interactions, UPESV can derive an effective video-based policy without the need for any other supervision.
@@
  • UPESV trains a video labeling model through three organically combined self-supervised tasks while imitating a policy by cloning the labeled videos. Each task performs its duties, and they together provide the video labeling model with comprehensive dynamics understanding and advanced action prediction.
@@
  • Extensive experiments in sixteen challenging procedurally generated environments demonstrate that UPESV exhibits superior sample efficiency of policy learning. It outperforms five current advanced baselines on 12/16 tasks with limited interactions.
@@ II-A Sample-efficient Policy Learning
Interacting with the environment to collect data for policy learning is a time-consuming and expensive process. In some special scenarios, such as autonomous driving [33] and real robot learning [34], the interactions can even lead to danger. To this end, many advanced methods are proposed to improve the RL sample efficiency. Model-based RL methods [9, 10, 35, 36] train extra world models that augment the RL experience, thereby improving sampling efficiency [37]. Self-supervised RL, i.e., RL auxiliary tasks [38, 39, 40, 41] alleviate the lack of supervision signals for representation learning in DRL, accelerating policy learning by improving the upstream representation understanding capabilities. At the same time, many researchers try to completely avoid accessing environments, training policies through imitation learning [19, 42] and offline RL [18, 17, 21]. However, the required expert supervision signals, including both expert reward and offline expert dataset, are also expensive and unavailable in many scenarios. Different from these expert supervisions, videos are easy to obtain from the internet currently. Our work aims to achieve completely unsupervised policy learning with the easily accessible action-free videos, in the absence of rewards and any other expert supervision.
Figure 2: UPESV framework. UPESV learns a video labeling model and a policy network jointly through three organically combined self-supervised tasks, where each is necessary and performs its own duties. The motivation and details of these three tasks (visual shift contrast, latent future reconstruction, and ground-truth action prediction) are provided separately in Sections III.A, III.B, and III.C. Simultaneously, we imitate a policy $\pi(a|o)$ by behavior cloning the labeled expert videos. The policy interacts with the reward-free environment, in turn enriching the self-supervised training data for the video labeling model. The labeling model and policy are improved iteratively, which we detail in Section III.D.
@@ II-B Learning Policies From Videos
Thanks to the improvement of the internet, it’s effortless to acquire massive and different kinds of videos currently. Many videos contain expert demonstrations in different fields, providing extensive knowledge for both humans and machines. However, due to the lack of action information, how to well utilize the videos poses a big challenge. One of the most direct ways is to improve existing strategy learning methods with the help of videos, such as video-based pre-training for RL [43, 44, 45, 46] and video-based intrinsic reward for RL [23]. They indeed improve the sample efficiency of policy learning but still require expert supervision (i.e., RL rewards). Imitation Learning from only Observations (ILO) [24, 28, 29] is proposed to completely decouple expert supervision from LFV. Most ILO methods try to extract expert rewards contained in the videos, employing RL to optimize the unsupervised reward expectation for policy imitation [24, 47, 48]. These inverse RL methods achieve considerable results on imitation performance but can’t take sample efficiency into account. At the same time, some researchers give up reward engineering for RL. By predicting expert actions from videos based on non-expert interactions [28, 49] or digging out expert dynamics hidden in the expert videos [29, 27], they achieve more efficient unsupervised policy learning without expert actions. Different from them, UPESV benefits from both non-expert interactions and expert videos by organically employing multiple self-supervised tasks, achieving efficient unsupervised policy cloning for complex video games.
@@ III Methodology
UPESV trains a policy network and a video labeling model jointly. The policy determines the action $a$ based on the observation $o$, defined as $\pi(a|o)$. The video labeling model predicts the real expert action between two neighboring expert observations in videos, formulated as an inverse dynamics model $V(a^{v}_{i}|o^{v}_{i},o^{v}_{i+1})$, where $a^{v}_{i}$ denotes the predicted action while $o^{v}_{i}$ and $o^{v}_{i+1}$ denote neighboring expert observations. $V$ consists of three networks: a feature encoder $f$, a latent predictor $g$, and an action projector $h$. $o^{v}_{i}$ and $o^{v}_{i+1}$ are separately encoded by $f$, concatenated, and processed by $g$ and $h$ to obtain the $a^{v}_{i}$, formulated as:
$a^{v}_{i}=h(g(f(o^{v}_{i}),f(o^{v}_{i+1}))).$ (1)
In practice, we also include a few available historical observations $o^{v}_{h}$ as additional input information. They are processed similarly to $o^{v}_{i}$ and we omit them in the remaining part for clarity. UPESV learns $V$ through three different but related self-supervised tasks over both expert videos $B^{v}$ and environmental interactions $B^{e}$, which we describe in the first three subsections. During the training of $V$, we simultaneously train the policy $\pi(a|o)$ by behavior cloning $V$-labeled expert videos, which we describe in Section III.D. The policy $\pi(a|o)$ collects self-supervised data for $V$, while the improved $V$ further enhances the policy $\pi(a|o)$. After a sample-efficient and iterative training process, we obtain the well-trained $V$ and $\pi$, as shown in Fig. 2.
@@ III-A1 Motivation
To infer actions from videos, the video labeling model $V(a^{v}_{i}|o^{v}_{i},o^{v}_{i+1})$ should be able to capture the difference between neighboring images. In most visual control domains, the relative difference is more related to decision-making than the absolute difference in visual inputs. Imagine two particles moving upward at different positions on the plane. They have different absolute positions, but the same relative position changes that correspond to their same actions. To this end, we want $V$ to focus more on relative position difference, which is achieved by driving it to ignore the absolute position difference. Therefore, a shift-matching task is employed, forcing the video labeling model to align two randomly shifted images with the same origin in the semantic space. This task enhances the visual understanding ability of $V$ in a target manner, which is necessary for $V$ to understand the environmental dynamics through another two tasks (detailed in the next two subsections) simultaneously.
@@ III-A2 Implementation
The visual shift contrast task forces the labeling model $V$ to align two randomly shifted images with the same origin in the latent semantic space. Concretely, a batch of observations $\{o^{v}_{i}\}_{i=1}^{N}$ is sampled from the expert video dataset $B^{v}$. Each observation is randomly shifted (up to $s$ units in any direction) twice to obtain two shifted images, $\hat{o}^{v}_{i}$ and $\check{o}^{v}_{i}$, where the two different hats upon ${o}^{v}_{i}$ denote two independent random shifts. They are separately encoded by $f$ and the momentum encoder $f^{\prime}$ to obtain two latent features, $f(\hat{o}^{v}_{i})$ and $f^{\prime}(\check{o}^{v}_{i})$. $f^{\prime}$ is not learned by gradient descent. It is updated through Exponential Moving Average (EMA) [50], similar to [38, 51]. Then, we define ($f(\hat{o}^{v}_{i})$,$f^{\prime}(\check{o}^{v}_{i})$) as positive pairs while ($f(\hat{o}^{v}_{i})$,$f^{\prime}(\check{o}^{v}_{j})$) as negative pairs. Contrastive learning [52] is conducted on the whole batch to bring the positive pair closer while pulling the negative pair further away from each other. A contrastive model that measures the distance is defined as follows:
$G_{c}(f(\hat{o}^{v}_{i}),f^{\prime}(\check{o}^{v}_{i}))=W\frac{u(f(\hat{o}^{v}_{i}))\cdot f^{\prime}(\check{o}^{v}_{i})}{||u(f(\hat{o}^{v}_{i}))||\cdot||f^{\prime}(\check{o}^{v}_{i})||},$ (2)
where $W$ is a trainable matrix. $u(\cdot)$ is a Multi-Layer Perceptron (MLP), which doesn’t change the dimension of the latent embedding. It is set to introduce asymmetry, thereby avoiding collapse to trivial solutions. Then, we define an InfoNCE loss [53] as the objective of visual shift contrast:
$\mathcal{L}_{VSC}=$ $\log G_{c}(f(\hat{o}^{v}_{i}),f^{\prime}(\check{o}^{v}_{i}))$ (3)
$-\log\frac{1}{N}\sum_{j=1}^{N}G_{c}(f(\hat{o}^{v}_{i}),f^{\prime}(\check{o}^{v}_{j})).$